OpenStudy (anonymous):

hey guys help me :D wuts the integral of \[e ^{-x ^{2}}\]?

OpenStudy (anonymous):

i got \[e ^{-x ^{2}}(-2x)\] but i don't think that's right...

OpenStudy (anonymous):

No, it requires the error function. It is \(\frac 1 2 \pi^{1/2} erf(x)\).

OpenStudy (anonymous):

wait wuts the error function? and how come the integration rules don't apply here????

OpenStudy (anonymous):

http://www.youtube.com/watch?v=FYNHt4AMxc0

OpenStudy (anonymous):

lol can't go on youtube right now :P

OpenStudy (anonymous):

doesn't it significantly change the integral value?

OpenStudy (anonymous):

Yes. It's essentially \(\sqrt \pi\).

OpenStudy (anonymous):

ok so lemme tell u the whole problem. im trying to find the area under the curve in the interval (0,0.922) for the function i posted, and im getting a negative value...the sq rt pi won't change the sign...the graph is above the x-axis so im suppose to get a positive value...

OpenStudy (anonymous):

It's positive.

OpenStudy (anonymous):